What is the density of water when:


The weight of an empty cylinder was 50 g, 20 mL of water was placed in the graduated cylinder, and the weight of the cylinder and water was 70 g.


1
Expert's answer
2021-11-24T12:41:20-0500

Weight of water=70-50=20g

Volume=20ml

"1ml=1g\/cm^3"

Therefore "20ml=20g\/cm^3"

"Density=\frac{mass}{volume}"

"Density=\frac{20g}{20g/cm^3}=1g/cm^3"


"1g\/cm^3=1000Kg\/M^3"


The density of water="1000Kg\/M^3"